OpenStudy (anonymous):

gcf factor. 5x^2+40x

OpenStudy (anonymous):

this question is really asking for the greatest common factor of the two terms. that will be \[5x\] since \[5x^2=5x\times x\] and \[40x=5x\times 8\]

OpenStudy (anonymous):

it also means you can factor as \[5x^2+40x=5x(x+8)\]

OpenStudy (anonymous):

however you might want to mention to your math teacher that there is no such thing as the greatest common factor of \[5x^2+40x\] only a greatest common factor of each term.

OpenStudy (anonymous):

greatest common factor= 5x because 5x(x+8)= 5x^2+40x

OpenStudy (anonymous):

5x =gcf 5x(x+8)

OpenStudy (anonymous):

How to factor d^2+7d+12=0 with two answers